CABOT CORPORATION Balance Sheet December 31, 2817 Assets Cash Short-term investments Accounts receivable, net Notes receivable (trade) Merchandise inventory Liabilities and Equity $12,e8 Accounts payable 28,888 Income taxes payable 38,15e Long-term note payable, secured by $15,58e 4,000 3,188 8, Accrued wages payable 5,580 65,488 mortgage on plant assets Prepaid expenses Plant assets, net Total assets 3,ee Common stock 154,3e8 Retained earnings 77,75e 249,758 Total liabilities and equity $ 249,75e These are short-term notes recelvable arising from customer (trade) sales. Requlrec Compute the following: (1) current ratio, (2) acid-test ratio, (3) days sales uncollected, (4) Inventory turnover, (5) days sales In Inventory, (6) debt-to-equity ratio, (7) times Interest earned, (8) profit margin ratio, (9) total asset turnover, (10) return on total assets, and (11) return on common stockholders equlty. (Do not round Intermedlete celculetions.) Complete this question by entering your answers in the tabs below Req 1 and 2 Req 3 Req 4 Req 5 Req 6 Req 7 Req 8 Req 9 Req 10 Req 11 Compute the debt-to-equity ratio Debt-to-Equity Ratio Choose Numerator: | 1 | Choose Denominator: | Debt-to-E quity Ratio = otal liabilities Total equity - Debt-to-equity ratio 2017: S 181.7501 0.21 to 1 < Req 5 Req 7>
Answer #1
(6) Debt-to-Equity Ratio
Choose Numerator: / Choose Denominator: = Debt-to-Equity Ratio
Total liabilities / Total equity = Debt-to-equity ratio
2017: 88000 / 161750 = 0.54 to 1

Working:

Total liabilities:
Accounts payable 15500
Accrued wages payable 4000
Income taxes payable 3100
Long-term note payable 65400
Total liabilities 88000
Total equity:
Common stock 84000
Retained earnings 77750
Total equity 161750